Two semiconductor materials have exactly the same properties except that material A has a

bandgap energy of 1.0 eV and material B has a bandgap energy of 1.2 eV. Determine the ratio

of ni, of material A to that of material B for T = 300 K.


1
Expert's answer
2021-09-09T10:51:58-0400

We have to use the relation for the holes on the semiconductor as:

"\cfrac{n_{i,A}^2}{n_{i,B}^2}=\large { \cfrac{e^{-\frac{E_g^A}{kT} } }{e^{-\frac{E_g^B}{kT} } } } = e^{({\frac{E_g^B-E_g^A}{kT}})}
\\ \text{we use the value for k in eV/K and substitute to find the relation:}


\\ \cfrac{n_{i,A}^2}{n_{i,B}^2}=e^{\frac{(1.2-1.0)eV}{(8.617 \times 10^{-5}\,eV/K)(300\,K)} }=e^{7.7366}
\\ \implies \cfrac{n_{i,A}}{n_{i,B}}=e^{3.8683}=47.86"


After substitution we find that the relation "\cfrac{n_{i,A}}{n_{i,B}} \text { is equal to } 47.86".


Reference:

  • Sears, F. W., & Zemansky, M. W. (1973). University physics.